OpenStudy (anonymous):

f(x) = 4x + 5 g(x) = 7x - 2 Find: (f - g)(7)

OpenStudy (blurbendy):

Same thing! See what happens if you substitute 7 in for x for both equations

OpenStudy (anonymous):

would i muiltiply by 7 at the end or no?

OpenStudy (blurbendy):

nope

OpenStudy (blurbendy):

just subtract what you get for both functions are putting 7 in

OpenStudy (anonymous):

thanks (:

OpenStudy (blurbendy):

-are +after

OpenStudy (anonymous):

what?

OpenStudy (blurbendy):

i said are instead of ater

OpenStudy (blurbendy):

after*

OpenStudy (anonymous):

whyy?

OpenStudy (anonymous):

can you help me out with this one @blurbendy f(x) = 4x + 5 h(x) = √5x - 4 Find: (hf)(4)

OpenStudy (blurbendy):

substitute 4 in for x for both equations then multiply is for h(x) is everything under the square root symbol or just 5x?

OpenStudy (anonymous):

everything.

OpenStudy (anonymous):

thats what i did but i didnt get any of the choice

OpenStudy (blurbendy):

sqrt[5(4) -4) = sqrt[16] = 4 f(4) = 4(4) + 5 = 21 21*4 = 84

OpenStudy (anonymous):

thanks (:

OpenStudy (blurbendy):

ur welcome

OpenStudy (anonymous):

im confused on this one, what do i substitiute? . f(x) = 2x - 7 g(x) = -4x2 - 6x Find: g(f(x))

OpenStudy (blurbendy):

so now we have to put f(x) into g(x) so, wherever you see an x in g(x), put f(x) in there g(2x-7) = -4(2x-7)^2 - 6(2x-7) (2x-7)^2 = (2x -7)(2x-7) = 4x^2 -14x - 14x + 49 = 4x^2 -28x + 49 -4 ( 4x^2 - 28x +49) = -16x^2 + 112x - 196 - 6(2x - 7) = -16x^2 + 112x - 196 - 12x + 42 = -16x^2 + 100x -154

OpenStudy (anonymous):

thank youuu (:

OpenStudy (blurbendy):

yw

OpenStudy (anonymous):

would this be the same? g(x) = 5x^2 - 4x h(x) = 3x + 9 Find: g(h(x))

OpenStudy (anonymous):

@blurbendy

OpenStudy (blurbendy):

hey i had to grab a snack. yes, the same idea applies